OpenStudy (anonymous):

Solve the equation or inequality. Check your solutions. (p/p-2)=(2/5) I NEED HELP!!!

OpenStudy (anonymous):

\[\frac{ p }{ p-2 } = \frac{ 2 }{ 5 }\] right?

OpenStudy (anonymous):

so you basically cross-multiply to get 2(p-2) = 5p so you open the brackets and solve for p and to check all you have to do is to replace the p's with -4/3 and see if you get 2/5 :)